OpenStudy (anonymous):

A true-false test consists of 20 questions, each of which has one correct answer: true, or false. One point is awarded for every correct answer, but one point is taken off for each wrong answer. Suppose a student answers every question by guessing at random, independently of other questions. Let S be the student’s score on the test. 1-Find E(S) 2-Find SE(S) 3-Find (no approximations, please) P(S=0)

OpenStudy (anonymous):

E(S)=0

OpenStudy (anonymous):

\[P(S=0)=\frac{1}{2^{10}}\]

OpenStudy (anonymous):

thank you! i needed it for homework!

OpenStudy (anonymous):

THANK YOU SOOO MUCH! The lecturer was rambling on with a prewritten script, and I couldn't understand most of what she was saying!

OpenStudy (anonymous):

the standard error score cannot be calculated like this

OpenStudy (kropot72):

@lucky12345 I agree. Therefore my attempt has been deleted.

OpenStudy (anonymous):

@kropot72 ans is wrong plzzzz help

OpenStudy (anonymous):

Find SE(S) ? help please

OpenStudy (anonymous):

SE(S)= 4.472 E(S-E(S))= 20-0 =20 E(S-E(S)) POWER 2 = 20 POWER 2 = 400 =400*0.05 =20 =SQRT OF 20 =4.472

OpenStudy (anonymous):

P(S=0) ans is 0.1761

OpenStudy (anonymous):

Exercise Set 6 Solution: http://fe9c2a32.linkbucks.com

OpenStudy (anonymous):

Exercise Set 7 Solution: http://f45b5978.linkbucks.com

OpenStudy (anonymous):

Exercise set 5 answers: http://4c51a450.linkbucks.com